# Enhanced detection of time-dependent dielectric structure: Rayleigh limit and quantum vacuum

### v3

@@ IV.2 Vacuum response
Eqs. (54, 55) imply for the vacuum response
$\widetilde{I}[\omega,\mathbf{r};|0\rangle\langle 0|]=\langle 0|A^{{\rm[s]}\dagger}_{y}[\omega,\mathbf{r}]A^{{\rm[s]}}_{y}[\omega,\mathbf{r}]|0\rangle=\frac{\gamma^{2}}{4\pi^{2}v^{2}}\times$
$\int{\rm d}\mathbf{q}\,q^{-1}(q-vq_{x})^{2}((q-vq_{x})^{2}-[1-v^{2}]q_{y}^{2})\times$
$\Big{|}{\cal V}[-\mathbf{q},-\gamma(q-vq_{x}),\omega+q,\mathbf{r}]\Big{|}^{2},$ (58)
@@
To simplify (58), we shall work with even a simpler model of a thin road
@@
We find from (58, 59) after going to spherical coordinates and changing variables:
$\widetilde{I}[\omega,\rho;|0\rangle\langle 0|]=\frac{\omega^{6}\gamma^{2}}{v^{2}(2\pi)^{3}}\int_{0}^{\infty}{\rm d}q\,q^{5}\int_{-1}^{1}{\rm d}\xi(1-v\xi)^{-2}\times$
$\Big{|}\widetilde{\varepsilon}\Big{[}\gamma q\omega;\frac{\omega(1-v\xi+q)}{v\gamma(1-v\xi)}\Big{]}\Big{|}^{2}\times$
[…]
where $K_{0}$ is Bessel’s $K_{0}$-function [31]. Its argument $\propto\sqrt{[1+q]^{2}-v^{2}}$ is non-negative. Since $\widetilde{I}$ in (IV.2) depends on $\rho$ only, it refers to cylindrical fields. This is due to the thin-road model (59) and the anisotropy introduced by the motion along the $x$-axis. The spherical far-field limit (23) does not apply to the moving dielectric e.g. because $|\widetilde{\varepsilon}|^{2}$ in (IV.2) does not depend on $x$, and hence does not nullify for $x\to\pm\infty$. The known asymptotics
@@
IV.3 Example
Let us focus on the following natural example in (IV.2). We assume in (59) a fixed linear size $2a$; cf. (27)  [^1] The sharp boundary assumption made in (62) will have to be revised, if more general quantities in (48) (which involve space-derivatives of $\widetilde{\varepsilon}$) are to be calculated. :
$\hat{\varepsilon}[\hat{\omega},x]=\theta(a-|x|)\,\hat{\varepsilon}[\hat{\omega}],$ (62)
where $\theta(x)$ is the step-function. Eq. (IV.2) now reads
@@ IV.3 Example
$\widetilde{I}[\omega,\rho;|0\rangle\langle 0|]=\frac{\omega^{4}\gamma^{4}}{8\pi^{5}}\int_{0}^{\infty}{\rm d}q\,q^{5}\int_{-1}^{1}{\rm d}\xi(1-v\xi+q)^{-2}\times$
$|\hat{\varepsilon}[\gamma q\omega]\,|^{2}\,K^{2}_{0}\left(\frac{\rho\omega}{v}\sqrt{[1+q]^{2}-v^{2}}\,\right)\times$ (63)
[…]
$\Omega\equiv\frac{\omega(1-v\xi+q)}{v\gamma(1-v\xi)}>0.$ (65)
In the ultra-relativistic limit $v\to 1$, $\gamma\to\infty$, we can replace $\sin^{2}[\Omega a]$ in (64) by $\frac{1}{2}$ and employ for the second term in (64):
$(1-v^{2})\int_{-1}^{1}{\rm d}\xi\,\frac{(1-\xi^{2})}{(1-v\xi+q)^{2}(1-v\xi)^{2}}\underset{v\to 1}{\longmapsto}0.$ (66)
@@
For $|\hat{\varepsilon}(\gamma q\omega)|^{2}$ in (63) we note the following relation that is valid for all materials in the considered large-frequency limit [26]:
$|\hat{\varepsilon}[\gamma q\omega]\,|^{2}\underset{\gamma\to\infty}{\longmapsto}\omega_{\rm e}^{4}(\gamma\omega q)^{-4},~{}~{}\omega_{\rm e}=\sqrt{4\pi N_{\rm e}{\rm e}^{2}/m_{\rm e}},$ (67)
@@
where $N_{\rm e}$ is the number of electrons in the sample, ${\rm e}$ and $m_{\rm e}$ are the electron charge and mass. The origin of (67) is that in the high-frequency limit, the material reacts to the electric field via its (free) electrons. Note that (67) is naturally incorporated into the Debye model.
Collecting (6467), we simplify (63) as:
$\widetilde{I}_{\rm ultra}[\omega,\rho;|0\rangle\langle 0|]=\frac{\omega_{\rm e}^{4}}{8\pi^{5}}\times$ (68)
$\int_{0}^{\infty}\frac{{\rm d}q\,K^{2}_{0}\left(\rho\omega\sqrt{[1+q]^{2}-1}\,\right)}{2+q}.$ (69)
Eq. (68) shows that the vacuum intensity in the ultra-relativistic limit does not directly depend on the linear size $L$. An indirect size dependence via $N_{\rm e}$ in (67) is there.
For $\rho\omega\gg 1$ (more precisely $3\lesssim\rho\omega$) the integral (69) is well approximated as $\frac{1}{4}(\rho\omega)^{-2}$, i.e.
$\widetilde{I}_{\rm ultra}[\omega,\rho;|0\rangle\langle 0|]=\frac{\omega_{\rm e}^{4}}{32\pi^{5}\rho^{2}\omega^{2}},\quad\rho\omega\gg 1.$ (70)
Eq. (70) shows that $\widetilde{I}_{\rm ultra}$ is a long-range effect, though it is not a far-field effect, because for cylindrical waves the far-field means intensity of order $\rho^{-1}$. For $\rho\omega\ll 1$, the integral in (69) is expressed by polylog functions and approximates as $\zeta\Big{[}\ln(\frac{1}{\rho\omega})\Big{]}^{3}$, where (as we found numerically) $0.280.37>\zeta>$ 0.28 for $\rho\omega<0.1$.
@@
Finally, let us return to (63, 64, 61) and note that out of the ultra-relativistic limit, the behavior of $\widetilde{I}[\omega,\rho;|0\rangle\langle 0|]$ in (63, 64) is (roughly) exponential for $\rho\omega$:
$\widetilde{I}[\omega,\rho;|0\rangle\langle 0|]={\cal F}(\rho\omega)\,e^{-2\rho\omega\sqrt{1-v^{2}}},$ (71)
@@
where ${\cal F}(\rho\omega)$ is not an exponential function of its argument. Eq. (71) is found from using (61) in (63, 64), changing the variable $w=\sqrt{(1+q)^{2}-v^{2}}$, and noting that $w\in[\sqrt{1-v^{2}},\infty]$.
@@ V Summary
The question asked in this paper is whether rectilinear motion, or its mimicking via space-time modulation of dielectric susceptibility, can lead to new mechanisms of dielectric structure determination. The question is dual to the Doppler metrology, where one employs scattered (or emitted) waves for measuring the velocity of a moving object [9]. Here we purposefully put the object into motion for determining its dielectric features. As a result of answering the above question, we uncovered several effects that may be interesting on their own.
We show that shining a light on the moving dielectric is not necessary, because there is a quantum vacuum response of a photodetector to a rectilinear moving, or space-time modulated (resting) dielectric. For the moving case this is a near-field effect (for cylindrical waves), but it can be long-range if the motion is ultra-relativistic. For the space-time modulated situation the vacuum effect pertains to far-field spherical waves. The latter (but not the former) effect has interesting—though certainly limited—analogies with the anomalous Doppler effect.
The scattering of single photons over a space-time modulated dielectric was also studied. Here we show explicitly that in the far-field zone of weak scattering the Rayleigh limit for structure determination (along the modulation axis) can be broken, i.e. a better resolution of the internal dielectric structure is possible.
@@ Appendix B Relations with Doppler physics
$\hat{\varepsilon}[k_{x},k_{y},k_{z}]$ is weakly dependent on $k_{y}$ and $k_{z}$ (i.e. the moving dielectric is a thin road).
@@ Appendix C Evaluation of vacuum contribution
where the function $\zeta$ has the following form in the far field [cf. (22, 23)]
@@
For a large $A$ (i.e., $A[\frac{\omega}{u}+\frac{1}{a}]\gg 1$) we can approximate $\delta(y)\simeq Ae^{-2A|y|}$. This leads to (30). Finally, (88) evaluates to (31, 32, 33).
@@
$r$ $=100\text{ cm};$ (98)
[…]
$\hbar$ $=1.054\cdot 10^{-27}\text{ erg}\cdot\text{s}.$ (100)
Eq. (32) equals $1.888\cdot 10^{-6}$ for the parameter values given in (93-100). Hence, (31) is $1.5858\cdot 10^{-7}$ in units of $\frac{\hbar\omega^{2}}{c^{3}}$. Note that the linear dimension of the sample $a=0.09\,\mu$m is smaller than the scattered field wavelength $\lambda\sim 6\,\mu$m.
@@ C.1.1 Single photon
Hence the photodetection signal from the modulated meta-material can be larger than the single-photon signal for a realistic range of parameters.
@@ C.1.2 Coherent state of a laser field
Let us now compare the estimates (106, 107) with the photodetection signal from a coherent state of the field. This corresponds to the laser light. Using the continuous-mode representation of the coherent state
@@
Summing up, the various contribution to the intensity (spectral density) at the frequency $\omega=3\cdot 10^{14}\text{ Hz}$ in units of $\frac{\hbar\omega^{2}}{c^{3}}$ are as follows [cf. (106, 107)]:
@@ Appendix D Minkowski’s formulation of relativistic continuous medium electrodynamics
where (117, 118, 119) is obtained from (116) upon using standard identities for the Levi-Civita tensor [26]:
@@
We assume a dielectrical situation $\mu=1$ in (124, 127), and find (5).

### v4

@@ IV.2 Vacuum response
Eqs. (56, 57) imply for the vacuum response
$\widetilde{I}[\omega,\mathbf{r};$ $|0\rangle\langle 0|]=\langle 0|A^{{\rm[s]}\dagger}_{y}[\omega,\mathbf{r}]A^{{\rm[s]}}_{y}[\omega,\mathbf{r}]|0\rangle=\frac{\gamma^{2}}{4\pi^{2}v^{2}}$
×$\times\int{\rm d}\mathbf{q}\,q^{-1}(q-vq_{x})^{2}((q-vq_{x})^{2}-[1-v^{2}]q_{y}^{2})$
×$\times\Big{|}{\cal V}[-\mathbf{q},-\gamma(q-vq_{x}),\omega+q,\mathbf{r}]\Big{|}^{2},$ (59)
@@
To simplify (59), we shall work with a simpler model of a thin rod (cf.  Sec. III.2)
@@
We now employ (60) in (58) that leads to simplifying (57) due to $\delta(\boldsymbol{\rho})$. Now in (59) we go to spherical coordinates, i.e., replace $(q_{x},q_{y},q_{z})$ by $q_{x}=q\cos\theta$, $q_{y}=q\sin\theta\cos\phi$, $q_{z}=q\sin\theta\sin\phi$, where $q=\sqrt{q_{x}^{2}+q_{y}^{2}+q_{z}^{2}}\geq 0$, $\theta\in[0,\pi]$, and $\phi\in[0,2\pi]$. Then, we take the integral over $\phi$, and introduce new integration variables $\xi=\cos\theta$ and $q^{\prime}=q(1-v\xi)/\omega$. Eventually, we find (denoting $q^{\prime}$ by $q$ again):
$\widetilde{I}[\omega,\rho;|0\rangle\langle 0|]$ $=\frac{\omega^{6}\gamma^{2}}{v^{2}(2\pi)^{3}}\int_{0}^{\infty}{\rm d}q\,q^{5}\int_{-1}^{1}{\rm d}\xi(1-v\xi)^{-2}$
×$\times\left|\widetilde{\varepsilon}\Big{[}\gamma q\omega;\frac{\omega(1-v\xi+q)}{v\gamma(1-v\xi)}\Big{]}\right|^{2}$
×[ $\times\left[\,1-\frac{(1-v^{2})(1-\xi^{2})}{2(1-v\xi)^{2}}\,\right]$
×$\times K^{2}_{0}\left(\frac{\rho\omega}{v}\sqrt{[1+q]^{2}-v^{2}}\,\right),$ (61)
where $\widetilde{\varepsilon}\Big{[}\gamma q\omega;\frac{\omega(1-v\xi+q)}{v\gamma(1-v\xi)}\Big{]}$ is defined via (58), where $(y,z)$ is omitted, and where $K_{0}$ is Bessel’s $K_{0}$-function [32]. Its argument $\frac{\rho\omega}{v}\sqrt{[1+q]^{2}-v^{2}}$ is non-negative. Since $\widetilde{I}$ in (IV.2) depends on $\rho$ only, it refers to cylindrical fields. This is due to the thin-rod model (60) and the anisotropy introduced by the motion along the $x$-axis. The spherical far-field limit (24) does not apply to the moving dielectric e.g. because $|\widetilde{\varepsilon}|^{2}$ in (IV.2) does not depend on $x$, and hence does not nullify for $x\to\pm\infty$. The known asymptotics
@@
IV.3 Example: point-like dielectric
We focus on the following simple example in (IV.2, 60):
$\hat{\varepsilon}[\hat{\omega},x]=\lambda\delta(x)\,\hat{\varepsilon}[\hat{\omega}],$ (63)
where $\lambda$ is a dimensional constant. Eqs. (60, 63) mean that we consider a point-like object.
Now the integral over $\xi$ in (IV.2) can be taken, and we get
@@ IV.3 Example: point-like dielectric
$\widetilde{I}[\omega,\rho;|0\rangle\langle 0|]=\frac{\omega^{6}\gamma^{4}\,\phi(v)\lambda^{2}}{(2\pi)^{3}}\int_{0}^{\infty}{\rm d}q\,q^{5}\left|\widetilde{\varepsilon}\Big{[}\gamma q\omega\Big{]}\right|^{2}$
×$\times K^{2}_{0}\left(\frac{\rho\omega}{v}\sqrt{[1+q]^{2}-v^{2}}\,\right),$ (64)
$\phi(v)=\frac{2}{v^{5}}(v-v^{3}(1-v^{2})-(1-v^{2})^{2}\,{\rm arctanh}[v]).~{}~{}~{}~{}~{}$ (65)
@@
We turn to the ultrarelativistic limit $v\to 1$ (i.e. $\gamma\to\infty$) in (64). For $|\hat{\varepsilon}(\gamma q\omega)|^{2}$ in (64) we note the following relation that is valid for all materials in the considered large-frequency limit [27]:
$|\hat{\varepsilon}[\gamma q\omega]\,|^{2}\underset{\gamma\to\infty}{\mapsto}\omega_{\rm e}^{4}(\gamma\omega q)^{-4},~{}~{}\omega_{\rm e}=\sqrt{4\pi N_{\rm e}{\rm e}^{2}/m_{\rm e}},$ (66)
@@
where $\omega_{e}$ is the plasma frequency with $N_{\rm e}$ being the electron density in the material; ${\rm e}$ and $m_{\rm e}$ are (resp.) the electron charge and mass. Note that $\gamma^{-4}$ in (66) matches to $\gamma^{4}$ in $\frac{\omega^{4}\gamma^{4}}{8\pi^{5}}$; see (64). The origin of (66) is that in the high-frequency limit, the material reacts to the electric field via its (free) electrons. Note that (66) is naturally incorporated into the Drude model.
We get from (64, 66):
$\widetilde{I}[\omega,\rho;|0\rangle\langle 0|]=\frac{2\omega^{2}\omega_{\rm e}^{2}\lambda^{2}}{(2\pi)^{3}}\,J(\rho\omega),$ (67)
$J(\rho\omega)=\int_{0}^{\infty}{\rm d}q\,qK^{2}_{0}\left(\rho\omega\sqrt{[1+q]^{2}-1}\,\right)$
$=\int_{0}^{\infty}{\rm d}u\,\frac{u(\sqrt{u^{2}+1}-1)}{\sqrt{u^{2}+1}}K^{2}_{0}\left(\rho\omega u\right).$ (68)
To present the asymptotic behavior of $J(\rho\omega)$ for $\rho\omega\gg 1$ and $\rho\omega\ll 1$ we write it as follows
$J(\rho\omega)=c_{1}(\rho\omega)(\rho\omega)^{-4},$ (69)
$J(\rho\omega)=c_{2}(\rho\omega)(\rho\omega)^{-2},$ (70)
where $c_{1}(\rho\omega)$ [$c_{2}(\rho\omega)$] is a slowly varying function for $\rho\omega\gg 1$ [$\rho\omega\ll 1$]. Now $c_{1}(1)=0.076$, $c_{1}(10)=0.163$ and $c_{1}(\rho\omega>20)\to 0.166$. Likewise, $c_{2}(\rho\omega)$ tends to $0.5$ for $\rho\omega\to 0$, e.g., $c_{1}(0.1)=0.35$, $c_{1}(0.01)=0.478$, etc.
Hence, (69) [(70)] describes $J(\rho\omega)$ at large [small] values of $\rho\omega$. The asymptotic behavior $(\rho\omega)^{-4}$ in (69) can be obtained by employing (62) in (68) and expanding $\frac{u(\sqrt{u^{2}+1}-1)}{\sqrt{u^{2}+1}}\simeq\frac{1}{2}u^{3}$ for small $u$, which is the dominant integration domain. Likewise, $(\rho\omega)^{-2}$ in (70) can be obtained by changing in (68) variables $u^{\prime}=\rho\omega u$, and noting that for a small $\rho\omega$ we have $\rho\omega\sim u^{\prime}$ in the dominant integration domain.
Eqs. (69, 70) show that in the ultrarelativistic limit, the detection of the vacuum signal is a long-range effect, though it is not a far-field effect, because for cylindrical waves the far-field means intensity $\propto\rho^{-1}$.
@@
Finally, let us return to (64, 62) and note that outside the ultra-relativistic limit, the behavior of $\widetilde{I}[\omega,\rho;|0\rangle\langle 0|]$ in (64) is (roughly) an exponential function of $\rho\omega$:
$\widetilde{I}[\omega,\rho;|0\rangle\langle 0|]={\cal F}(\rho\omega)\,e^{-2\frac{\rho\omega}{v}\sqrt{1-v^{2}}},$ (71)
@@
where ${\cal F}(\rho\omega)$ is not an exponential function of its argument. Eq. (71) is found from using (62) in (64), changing the variable $w=\sqrt{(1+q)^{2}-v^{2}}$, and noting that $w\in[\sqrt{1-v^{2}},\infty]$. Eq. (71) also describes the behavior of $\widetilde{I}[\omega,\rho;|0\rangle\langle 0|]$ for $v\to 0$.
@@ V Summary
The question asked in this paper is whether rectilinear motion, or space-time modulation of dielectric susceptibility, can lead to new mechanisms of dielectric structure determination. These two problems are studied together because the space-time modulation makes the dielectric susceptibility of a resting dielectric formally similar to that of a moving dielectric. The literature suggests that experimentally more feasible space-time modulation can model relativistically moving dielectrics [24]. Experimentally studying systems in relativistic motion is undoubtedly more challenging.
The question is dual to Doppler metrology, where one employs scattered (or emitted) waves to measure the velocity of a moving object [9]. Here we purposefully put the object into motion (or space-time modulate it) to determine its dielectric features. As a result of answering the above question, we have uncovered several effects that may be interesting on their own.
We show that radiating a light on the moving dielectric is not necessary because there is a quantum vacuum response of a photodetector to such a motion. This is a near-field effect (for cylindrical waves), but it can be long-range if the motion is ultra-relativistic; see section IV. The vacuum response is also present in the space-time modulated situation, but now the important point is that the response can be detected in the far-field zone (for spherical waves); cf. section III. This effect has interesting – though certainly limited – analogies with the anomalous Doppler effect; see section III.2.1.
We also studied the scattering of single photons over a space-time modulated dielectric. Here we show explicitly that in the far-field zone of weak scattering the Rayleigh limit for structure determination (along the modulation axis) can be violated, i.e. a better resolution of the internal dielectric structure is possible.
@@ Appendix B Relations with Doppler physics
$\hat{\varepsilon}[k_{x},k_{y},k_{z}]$ is weakly dependent on $k_{y}$ and $k_{z}$ (i.e. the moving dielectric is a thin rod).
@@ Appendix C Evaluation of vacuum contribution
where the function $\zeta$ has the following form in the far field [cf. (23, 24)]
@@
For a large $A$ (i.e., $A[\frac{\omega}{u}+\frac{1}{a}]\gg 1$) we can approximate $\delta(y)\simeq Ae^{-2A|y|}$. This leads to (31). Finally, (88) evaluates to (32, 33, 34). Fig. 3 presents a numerical illustration of (32) with parameters discussed below.
@@
$r$ $=100\text{ cm}.$ (98)
Eq. (33) equals $1.888\cdot 10^{-6}$ for the parameter values given in (93-98). Hence, (32) is $1.5858\cdot 10^{-7}$ in units of $\frac{\hbar\omega^{2}}{c^{3}}$; see Fig. 3 for illustration. Note that the linear dimension of the sample $a=0.09\,\mu$m is smaller than the scattered field wavelength $\lambda\sim 6\,\mu$m.
@@ C.1.1 Single photon
Hence the photodetection signal from the modulated meta-material can be larger than the single-photon signal for a realistic range of parameters. The behavior of (103) is illustrated in Fig. 4.
@@ C.1.2 Coherent state of a laser field
Let us now compare the estimates (104, 105) with the photodetection signal from a coherent state of the field. This corresponds to the laser light. Using the continuous-mode representation of the coherent state
@@
Summing up, the various contribution to the intensity (spectral density) at the frequency $\omega=3\cdot 10^{14}\text{ Hz}$ in units of $\frac{\hbar\omega^{2}}{c^{3}}$ are as follows [cf. (104, 105)]:
@@ Appendix D Minkowski’s formulation of relativistic continuous medium electrodynamics
where (115, 116, 117) is obtained from (114) upon using standard identities for the Levi-Civita tensor [27]:
@@
We assume a dielectrical situation $\mu=1$ in (122, 125), and find (5).